# Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments

### v1

@@ Appendix B Proofs
$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (33)
@@
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (34)
Since the bootstrap means converge to the empirical ones we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$. We also have, as in (10), where empirical means are replaced by bootstrap means,
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (35)
for any $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$, with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right)=O_{p}(1)$. As a result, we deduce from Lemma 2.4, inequality (34) and Lemma 2.9 in Newey and McFadden, (1994) that
@@
We conclude that $\boldsymbol{\theta}^{*}\to\boldsymbol{\theta}_{0}$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). We now prove that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have the same asymptotic distribution. By definition of $\widehat{\boldsymbol{\theta}}$ and Taylor expansion we have
$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (36)
@@
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (37)
where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (36) and (37), we deduce
@@
$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (38)
@@
$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (39)
@@
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (45)

### v3

@@ Appendix B Proofs
Proof.
of Lemma 5.1
Note that
$\widehat{\beta}_{1}=g(\widehat{\mu}_{X}^{1},\ldots,\widehat{\mu}_{X}^{K},\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K}),$
with $g:\mathbb{R}^{K+K}\to\mathbb{R}$ defined as follows,
$g(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$ $=\frac{\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}}{\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-\left(\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\right)^{2}},$
$=\frac{\mbox{Cov}_{w}(X,Y)}{\mbox{Var}_{w}(X)},$ (B.9)
with the notations $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$, $\mu_{Y,w}=\sum_{j=1}^{K}w_{j}\mu_{Y}^{j}$, $\mbox{Cov}_{w}(X,Y)=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-(\mu_{X,w})^{2}$. The gradient $\nabla g$ of $g$, evaluated at point $(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, is equal to
$\nabla g$ $=\begin{pmatrix}\frac{w_{1}(\mu_{Y}^{1}-\mu_{Y,w})}{\mbox{Var}_{w}(X)}-\frac{2w_{1}\left(\mu_{X}^{1}-\mu_{X,w}\right)\mbox{Cov}_{w}(X,Y)}{(\mbox{Var}_{w}(X))^{2}}\\ \vdots\\ \frac{w_{K}(\mu_{Y}^{K}-\mu_{Y,w})}{\mbox{Var}_{w}(X)}-\frac{2w_{K}\left(\mu_{X}^{K}-\mu_{X,w}\right)\mbox{Cov}_{w}(X,Y)}{(\mbox{Var}_{w}(X))^{2}}\\ \frac{w_{1}\left(\mu_{X}^{1}-\mu_{X,w}\right)}{\mbox{Var}_{w}(X)}\\ \vdots\\ \frac{w_{K}\left(\mu_{X}^{K}-\mu_{X,w}\right)}{\mbox{Var}_{w}(X)}\end{pmatrix}.$
As in the proof of Proposition 4.1, we get that $\sqrt{n}\left(\widehat{\beta}_{1}-\beta_{1}\right)\rightsquigarrow\mathcal{N}(0,\sigma^{2}_{\beta_{1}})$ with $\sigma^{2}_{\beta_{1}}=\nabla g^{T}\Gamma_{\mu}\nabla g$, so that
$\sigma^{2}_{\beta_{1}}$ $=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left[\sigma^{2}_{X,k}\left(\mu_{Y}^{k}-\mu_{Y,w}-2\beta_{1}\left(\mu_{X}^{k}-\mu_{X,w}\right)\right)^{2}+\sigma^{2}_{Y,k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\right]$
$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left[\sigma^{2}_{X,k}\left(-\beta_{1}\left(\mu_{X}^{k}-\mu_{X,w}\right)\right)^{2}+\sigma^{2}_{Y,k}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\right]$
$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)$ (B.10)
remarking that $\beta_{1}=\mbox{Cov}_{w}(X,Y)/\mbox{Var}_{w}(X,Y)$, $\beta_{0}=\mu_{Y,w}-\beta_{1}\mu_{X,w}$ as well as $\beta_{0}=\mu_{Y}^{k}-\beta_{1}\mu_{X}^{k}$.
@@
$\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\mu_{Y}^{k,*}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*})^{2}+\left(\sigma_{Y,k}^{*}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (B.11)
@@
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|$ $\leq\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|+\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\right|.$ (B.12)
Since the bootstrap means converge to the empirical ones we deduce with the continuous mapping theorem that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ in probability, when $n_{\min}$ tends to infinity, so that $\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\to\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$. We also have, as in (B.3), where empirical means are replaced by bootstrap means,
$\left|\varphi_{n}^{*}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}^{*}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right),$ (B.13)
for any $(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\in\Theta$, with $\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k,*}+A_{n}^{k,*}\right)=O_{p}(1)$. As a result, we deduce from Lemma 4.2, inequality (B.12) and Lemma 2.9 in Newey and McFadden, (1994) that
@@
We conclude that $\boldsymbol{\theta}^{*}\to\boldsymbol{\theta}_{0}$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994).
We now prove that $\sqrt{n}\left(\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}\right)$ and $\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)$ have the same asymptotic distribution. By definition of $\widehat{\boldsymbol{\theta}}$ and Taylor expansion we have
$\nabla\varphi_{n}(\widehat{\boldsymbol{\theta}})=\nabla\varphi_{n}(\boldsymbol{\theta}_{0})+\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0}\right)=0$ (B.14)
@@
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}^{*})=\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\left(\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{0}\right)=0$ (B.15)
where $\overline{\boldsymbol{\theta}}^{*}$ belongs, componentwise, to the segment between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}^{*}$. Combining (B.14) and (B.15), we deduce
@@
$=\left(\left(\nabla_{00}^{*}\varphi_{n}(\overline{\boldsymbol{\theta}}^{*})\right)^{-1}-\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\right)\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})+\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)$ (B.16)
@@
$\boldsymbol{\theta}^{*}-\widehat{\boldsymbol{\theta}}$ $=\left(\nabla_{00}\varphi_{n}(\overline{\boldsymbol{\theta}})\right)^{-1}\left(\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})\right)+o_{P}(n^{-1/2}).$ (B.17)
@@
$\nabla\varphi_{n}^{*}(\boldsymbol{\theta}_{0})-\nabla\varphi_{n}(\boldsymbol{\theta}_{0})=\footnotesize\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left((\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k,*})-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}-\boldsymbol{\mu}_{X}^{k,*}\right)\right)\\ \begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\mu_{Y}^{k,*}-\mu_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\mu}_{X}^{k,*}\right)\boldsymbol{\mu}_{X}^{k,*}+\left(\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}\right]\end{pmatrix}\\ \sum_{k=1}^{K}\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-\frac{\sigma_{Y,k}^{*}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\boldsymbol{\Gamma}_{X}^{k,*}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix},$ (B.18)
@@
Appendix C Supplementary figures
Figure C.1: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination. The method labels on the left: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.